# Search for the charged lepton flavor violating decay $\psi(3686)\to e^{\pm}\mu^{\mp}$

### v1

@@ Abstract
In the Standard Model (SM), quarks change flavor through electroweak transitions, with rates codified by the Cabibbo-Kobayashi-Maskawa matrix. Similarly, neutrinos oscillate between electron, muon, and tau lepton flavors according to rates given by the Pontecorvo-Maki-Nakagawa-Sakata matrix. Lepton flavor violation is therefore an established fact, but only for neutral neutrinos ref::neutrino_1 ; ref::neutrino_2 . The discovery of non-zero neutrino masses and neutrino oscillations guarantees that charged lepton flavor violation (CLFV) can occur in SM loops, but with very small branching fractions (BFs), which are experimentally undetectable. Therefore, any detection of CLFV would be an unambiguous signal of new physics beyond the SM ref::clfvreview_1 .
@@
Over the past several decades, searches for CLFV processes have been performed in the lepton, pseudo-scalar meson, vector meson, and boson sectors ref::CLFVreview . For example, in the lepton sector, searches for CLFV have been carried out in muon and tau decays, such as $\mu^{+}\to e^{+}\gamma$ ref::mu2egamma , $\mu^{+}\to e^{+}e^{-}e^{+}$ ref:mu2eee , $\mu^{-}N\to e^{-}N$ ref::muN2eN_1 ; ref::muN2eN_2 , $\mu^{-}N\to e^{-}N^{\prime}$ ref::muN2eNprime , $\tau^{\pm}\to l^{\pm}\gamma$ ref::tau2lgamma , $\tau\to lll$ ref::tau2lll_1 ; ref::tau2lll_2 , and $\tau\to hl$ ref::tau2hl_1 ; ref::tau2hl_2 ; ref::tau2hl_3 (where $l$ denotes a lepton and $h$ denotes a hadron, such as the $\pi^{0}$, $\eta^{(\prime)}$, $\phi$, $\omega$, $K^{*0}$, or $\rho^{0}$). However, no evidence of CLFV has yet been found. In the quarkonium sector, several CLFV decays of the $J/\psi$ and $\Upsilon(nS)$ have been investigated by the BESIII, BaBar, and Belle experiments, and no evidence for a signal has been found. Upper limits (UL) on the BFs of the decays $J/\psi\to e\mu$, $J/\psi\to e\tau$, $\Upsilon(1S)\to e\mu$, $\Upsilon(1S)\to\mu\tau$, $\Upsilon(1S)\to e\tau$, and $\Upsilon(3S)\to e\mu$ are currently $4.5\times 10^{-9}$ ref::jpsi2emu , $7.5\times 10^{-8}$ ref::jpsi2etau , $3.6\times 10^{-7}$ ref::upsilon2ll_1 , $2.6\times 10^{-6}$ ref::upsilon2ll_1 , $2.4\times 10^{-6}$ ref::upsilon2ll_1 , and $3.6\times 10^{-7}$ ref::upsilon2ll_2 at $90\%$ confidence level (C.L.), respectively.
@@
Theoretically, CLFV can be described by various new physics models, such as the two-Higgs doublet model with extra Yukawa couplings ref::yukawa , a model that constitutes a simple example of tree-level off-diagonal Majorana couplings not suppressed by neutrino masses ref::majorana , and a series of models based on super-symmetry (SUSY) ref::susy_gut ; ref::susy_leptons ; ref::susy_neutrino ; ref::susy_zprime . For the CLFV decays of charmonium, models involving SUSY-based grand unified theories ref::clfvpsip_gut , SUSY ref::clfvpsip_susy , and TC2 ref::clfvpsip_tc2 with various new bosons (scalars, vectors), predict their decay rates at the order of $10^{-8}$ to $10^{-16}$ ref::theory_clfvpsip_1 ; ref::theory_clfvpsip_2 ; ref::theory_clfvpsip_3 ; ref::theory_clfvpsip_4 ; ref::theory_clfvpsip_5 ; ref::theory_clfvpsip_6 .
@@
The BESIII analysis reported in this Letter is the first search for CLFV in the decay $\psi(3686)\to e^{\pm}\mu^{\mp}$. In this Letter, $(2367.0\pm 11.1)\times 10^{6}$ $\psi(3686)$ events collected in 2009 and 2021 at $\sqrt{s}=3.686~{}\rm GeV$ ref:numpsip are analyzed. The $e^{+}e^{-}$ collision events collected at $\sqrt{s}=3.773$, $3.65$, and $3.682~{}\rm GeV$ ref:data-3650-3773 ; ref:data-3773 , corresponding to the integrated luminosities of $7926.8$, $454.33$, and $404~{}\rm pb^{-1}$, respectively, are used to estimate the continuum background.
@@
The BESIII detector Ablikim:2009aa records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring Yu:IPAC2016-TUYA01 in the center-of-mass (c.m.) energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region. The data is monitored in real time using the Bhabha process and the Data Quality Monitoring system EcmsMea . The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate muon counters (MUC) identification modules interleaved with steel. The charged-particle momentum resolution at $1~{}{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the plastic scintillator TOF barrel region is 68 ps, while that in the end cap region was upgraded in 2015 from 110 ps to 60 ps using multi-gap resistive plate chamber technology ref:tofupgrade . About 95% of the dataset used in this analysis benefits from this upgrade.
@@
Monte Carlo (MC) simulated data samples produced with a geant4-based geant4 software package, which includes the geometric description Huang:2022wuo of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc ref:kkmc . The inclusive MC sample includes the production of the $\psi(3686)$ resonance, the ISR production of the $J/\psi$, and the continuum processes incorporated in kkmc ref:kkmc . All particle decays are modelled with evtgen ref:evtgen using BFs either taken from the Particle Data Group ref:pdg , when available, or otherwise estimated with lundcharm ref:lundcharm . Final state radiation from charged final state particles is incorporated using the photos package photos2 . Samples of $1\times 10^{5}$ events are generated separately for the experimental conditions of 2009 and 2021 to study the selection criteria. Approximately $2.406\times 10^{9}$ $\psi(3686)$ events from the inclusive MC sample are used to estimate the backgrounds from $\psi(3686)$ decays. To improve the statistics of the MC sample for background studies, an equivalent of $2.406\times 10^{10}$ $\psi(3686)$ exclusive events are generated, including decay channels such as $\psi(3686)\to e^{+}e^{-},~{}\mu^{+}\mu^{-},~{}\tau^{+}\tau^{-},~{}K^{+}K^{-}$, and $p\bar{p}$. To avoid involuntary bias, a semi-blind analysis method is employed. The full dataset is analyzed after the analysis procedure is finalized by using the MC simulation sample and $10\%$ of the full data sample.
@@
Figure 1: Distributions of $E/p$ (a) and $\chi^{e}_{dE/dx}$ (b) for $e$, $\mu$, $\pi$, $K$, and $p$ samples from MC simulations of $\psi(3686)\to e^{+}e^{-}$, $\mu^{+}\mu^{-}$, $\pi^{+}\pi^{-}$, $K^{+}K^{-}$, and $p\bar{p}$, respectively. In (a), the blue arrow indicates $E/p>0.94c$ for electron selection. In (b), the red arrows indicate $-1.2<\chi^{e}_{dE/dx}<1.2$ for electron selection, and the blue arrow indicates $\chi^{e}_{dE/dx}<-1.1$ for muon selection.
@@
Strict particle identification (PID) for the electron and muon is based on information provided by the MDC, EMC and MUC sub-detectors. For electron PID, the energy-to-momentum ratio $E/p$ is required to be greater than $0.94c$, as shown in Fig. 1(a), where $E$ is the deposited energy in the EMC and $p$ is the magnitude of the momentum measured by the MDC. In addition, $\chi^{e}_{dE/dx}$ is required to satisfy $|\chi^{e}_{dE/dx}|<1.2$, as shown in Fig. 1(b), where $\chi^{e}_{dE/dx}$ is the difference between the measured and expected specific ionization energy loss $dE/dx$ under the electron hypothesis, normalized by the $dE/dx$ resolution. Furthermore, since electrons cannot pass through the EMC and reach the MUC, the maximum number of hits in any layer of the MUC for the electron is required to be zero. For muon PID, an energy window of $[0.1,~{}0.3]~{}\rm GeV$ is applied, where the energy is the deposited energy measured by the EMC. In order to lower the probability to misidentify pions or kaons as muons, the penetration depth of the muon in the MUC is required to be greater than $20~{}\rm cm$, as shown in Fig. 2(a). To obtain a well-reconstructed MUC track, the number of layers penetrated by the muon in the MUC must be greater than three, and the fitted $\chi^{2}_{\rm MUC}$ for the muon in the MUC must be less than 100, as shown in Fig. 2(b). Additionally, to suppress misidentification due to electrons, the value of $\chi^{e}_{dE/dx}$ for the muon is required to be less than $-1.1$, as shown in Fig. 1(b).
@@
Figure 2: Distributions of the MUC depth (a) and $\chi^{2}_{\rm MUC}$ (b) for different particle samples from MC simulations. The blue arrows indicate the muon selection criteria: MUC depth $>20~{}\rm cm$ in (a) and $\chi^{2}_{\rm MUC}<100$ in (b).
@@
To suppress the background from Bhabha scattering events, in which one of the electrons is misidentified as a muon, additional selections are applied to the muon candidates. Muon candidates with an MDC track polar angle satisfying $0.860.82<|\!\cos\theta|<$ 0.86 are vetoed because of the gap of the EMC. The track direction of the muon candidate in the MDC and MUC is constrained. The polar angle difference between the track in the MUC and MDC is required to be less than $20^{\circ}$, and the azimuthal angle difference between the track in the MUC and MDC is required to be less than $30^{\circ}$. A detailed discussion can be found in Ref. ref:visualbes3 .
@@
Since the final state electron and muon are back-to-back, the opening angles in the polar and azimuthal directions are required to be less than $1.4^{\circ}$ and $1.7^{\circ}$, respectively. The opening angle in the polar direction is defined as $|\Delta\theta|=|180^{\circ}-(\theta_{1}+\theta_{2})|$, where $\theta_{1}$ and $\theta_{2}$ are the corresponding polar angles of the electron and muon, respectively. The opening angle in the azimuthal direction is defined as $|\Delta\phi|=|180^{\circ}-|\phi_{1}-\phi_{2}||$, where $\phi_{1}$ and $\phi_{2}$ are the azimuthal angles of the electron and muon, respectively. All angles are evaluated in the c.m. system of the $e^{+}e^{-}$ collisions.
@@
To suppress background from the processes $e^{+}e^{-}\to\gamma e^{+}e^{-}$ and $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$, the number of detected photons is required to be zero The photon candidates used for this veto are reconstructed using showers in the EMC detector. The showers are required to have a deposited energy greater than $25~{}\rm MeV$ in the barrel region ($|\!\cos\theta|<0.80$) and $50~{}\rm MeV$ in the end-cap region ($0.920.86<|\!\cos\theta|<$ 0.92). The difference between the EMC time and the event start time for each photon must be within [0, 700] ns to suppress electronic noise and showers unrelated to the event. To exclude showers originating from the final state charged tracks, the opening angle between the position of each shower in the EMC and the closest extrapolated charged track must be greater than 10 degrees.
@@
The normalized momentum sum $|\sum_{i}\vec{p_{i}}|/\sqrt{s}$ and the normalized energy sum $\sum_{i}E_{i}/\sqrt{s}$ are used to determine the number of signal events in the full dataset, where $\vec{p}_{i}$ is the momentum and $E_{i}$ is the reconstructed energy of final state particle $i$, and $\sqrt{s}$ is the c.m. energy. The signal region is defined by $|\sum_{i}\vec{p_{i}}|/\sqrt{s}\leq 0.03$ and $0.97\leq\sum_{i}E_{i}/\sqrt{s}\leq 1.04$. The selection criteria discussed above are optimized by maximizing the figure-of-merit (FOM) suggested by Punzi ref:punzi , given by

### v2

@@ Abstract
keywords:
BESIII, $e^{+}e^{-}$ experiments, charmonium, lepton flavor violating decay
1 Introduction
@@
In the Standard Model (SM), quarks change flavor through electroweak transitions, with rates codified by the Cabibbo-Kobayashi-Maskawa matrix. Similarly, neutrinos oscillate between electron, muon, and tau lepton flavors according to rates given by the Pontecorvo-Maki-Nakagawa-Sakata matrix, which is beyond SM [1]. However, the lepton flavor violation has not been discovered in the sector of charged leptons. The discovery of non-zero neutrino masses and neutrino oscillations guarantees that charged lepton flavor violation (CLFV) can occur in SM loops, but with very small branching fractions (BFs), which are experimentally undetectable. Therefore, any detection of CLFV would be an unambiguous signal of new physics beyond the SM [2].
@@
Over the past several decades, searches for CLFV processes have been performed in the lepton, pseudo-scalar meson, vector meson, and boson sectors [3]. For example, in the lepton sector, searches for CLFV have been carried out in muon and tau decays, such as $\mu^{+}\to e^{+}\gamma$ [4], $\mu^{+}\to e^{+}e^{-}e^{+}$ [5], $\mu^{-}N\to e^{-}N$ [6, 7], $\mu^{-}N\to e^{-}N^{\prime}$ [8], $\tau^{\pm}\to l^{\pm}\gamma$ [9], $\tau\to lll$ [10, 11], and $\tau\to hl$ [12, 13] (where $l$ denotes a lepton and $h$ denotes a hadron, such as the $\pi^{0}$, $\eta^{(\prime)}$, $\phi$, $\omega$, $K^{*0}$, or $\rho^{0}$). However, no evidence of CLFV has yet been found. In the quarkonium sector, several CLFV decays of the $J/\psi$ and $\Upsilon(nS)$ have been investigated by the BESIII, BaBar, and Belle experiments, and no evidence for a signal has been found. Upper limits (UL) on the BFs of the decays $J/\psi\to e\mu$, $J/\psi\to e\tau$, $\Upsilon(1S)\to e\mu$, $\Upsilon(1S)\to\mu\tau$, $\Upsilon(1S)\to e\tau$, and $\Upsilon(3S)\to e\mu$ are currently $4.5\times 10^{-9}$ [14], $7.5\times 10^{-8}$ [15], $3.6\times 10^{-7}$, $2.6\times 10^{-6}$, $2.4\times 10^{-6}$ [16], and $3.6\times 10^{-7}$ [17] at $90\%$ confidence level (C.L.), respectively.
@@
Theoretically, CLFV can be described by various new physics models, such as the two-Higgs doublet model with extra Yukawa couplings [18], a model that constitutes a simple example of tree-level off-diagonal Majorana couplings not suppressed by neutrino masses [19], and a series of models based on super-symmetry (SUSY) [20, 21, 22, 23]. For the CLFV decays of charmonium, models involving SUSY-based grand unified theories [24], SUSY [25], and TC2 [26] with various new bosons (scalars, vectors), predict their decay rates at the order of $10^{-8}$ to $10^{-16}$ [27, 28, 29, 30, 31, 32].
@@
The BESIII analysis reported in this Letter is the first search for CLFV in the decay $\psi(3686)\to e^{\pm}\mu^{\mp}$. In this Letter, $(2367.0\pm 11.1)\times 10^{6}$ $\psi(3686)$ events collected in 2009 and 2021 at $\sqrt{s}=3.686~\rm GeV$ [33] are analyzed. The $e^{+}e^{-}$ collision events collected at $\sqrt{s}=3.773$, $3.65$, and $3.682~\rm GeV$ [34, 35, 36], corresponding to the integrated luminosities of $7926.8$, $454.33$, and $404~\rm pb^{-1}$, respectively, are used to estimate the continuum background.
2 BESIII Detector and data samples
@@
The BESIII detector [37] records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring [38] in the center-of-mass (c.m.) energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region. The data is monitored in real time using the Bhabha process and the Data Quality Monitoring system [39]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate muon counters (MUC) identification modules interleaved with steel. The charged-particle momentum resolution at $1~{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the plastic scintillator TOF barrel region is 68 ps, while that in the end cap region was upgraded in 2015 from 110 ps to 60 ps using multi-gap resistive plate chamber technology [40]. About 95% of the dataset used in this analysis benefits from this upgrade.
@@
Monte Carlo (MC) simulated data samples produced with a geant4-based [41] software package, which includes the geometric description [42, 43] of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [44, 45]. The inclusive MC sample includes the production of the $\psi(3686)$ resonance, the ISR production of the $J/\psi$, and the continuum processes incorporated in kkmc. All particle decays are modelled with evtgen [46, 47] using BFs either taken from the Particle Data Group [48], when available, or otherwise estimated with lundcharm [49, 50]. Final state radiation from charged final state particles is incorporated using the photos package [51]. Samples of $1\times 10^{5}$ events are generated separately for the experimental conditions of 2009 and 2021 to study the selection criteria. Approximately $2.406\times 10^{9}$ $\psi(3686)$ events from the inclusive MC sample are used to estimate the backgrounds from $\psi(3686)$ decays. To improve the statistics of the MC sample for background studies, an equivalent of $2.406\times 10^{10}$ $\psi(3686)$ exclusive events are generated, including decay channels such as $\psi(3686)\to e^{+}e^{-},~\mu^{+}\mu^{-},~\tau^{+}\tau^{-},~K^{+}K^{-}$, and $p\bar{p}$. To avoid involuntary bias, a semi-blind analysis method is employed. The full dataset is analyzed after the analysis procedure is finalized by using the MC simulation sample and $10\%$ of the full data sample.
3 Event Selection
@@
Figure 1: Distributions of $E/p$ (a) and $\chi^{e}_{dE/dx}$ (b) for $e$, $\mu$, $\pi$, $K$, and $p$ samples from MC simulations of $\psi(3686)\to e^{+}e^{-}$, $\mu^{+}\mu^{-}$, $\pi^{+}\pi^{-}$, $K^{+}K^{-}$, and $p\bar{p}$, respectively. In (a), the blue arrow indicates $E/p>0.94c$ for electron selection. In (b), the red arrows indicate $1.2<\chi^{e}_{dE/dx}<1.2$ for electron selection, and the blue arrow indicates $\chi^{e}_{dE/dx}<-1.1$ for muon selection.
@@
Strict particle identification (PID) for the electron and muon is based on information provided by the MDC, EMC and MUC sub-detectors. For electron PID, the energy-to-momentum ratio $E/p$ is required to be greater than $0.94c$, as shown in Fig. 1(a), where $E$ is the deposited energy in the EMC and $p$ is the magnitude of the momentum measured by the MDC. In addition, $\chi^{e}_{dE/dx}$ is required to satisfy $|\chi^{e}_{dE/dx}|<1.2$, as shown in Fig. 1(b), where $\chi^{e}_{dE/dx}$ is the difference between the measured and expected specific ionization energy loss $dE/dx$ under the electron hypothesis, normalized by the $dE/dx$ resolution. Furthermore, since electrons cannot pass through the EMC and reach the MUC, the maximum number of hits in any layer of the MUC for the electron is required to be zero. For muon PID, an energy window of $0.1,~0.3]~\rm GeV$ is applied, where the energy is the deposited energy measured by the EMC. In order to lower the probability to misidentify pions or kaons as muons, the penetration depth of the muon in the MUC is required to be greater than $20~\rm cm$, as shown in Fig. 2(a). To obtain a well-reconstructed MUC track, the number of layers penetrated by the muon in the MUC must be greater than three, and the fitted $\chi^{2}_{\rm MUC}$ for the muon in the MUC must be less than 100, as shown in Fig. 2(b). Additionally, to suppress misidentification due to electrons, the value of $\chi^{e}_{dE/dx}$ for the muon is required to be less than 1.1-1.1, as shown in Fig. 1(b).
@@
Figure 2: Distributions of the MUC depth (a) and $\chi^{2}_{\rm MUC}$ (b) for different particle samples from MC simulations. The blue arrows indicate the muon selection criteria: MUC depth $20~\rm cm$ in (a) and $\chi^{2}_{\rm MUC}<100$ in (b).
@@
The energy leakage at $|\cos\theta|=0$ causes the rising of the rate of the misidentification from electron to muon. To further suppress the background from Bhabha scattering events, in which one of the electrons is misidentified as a muon, additional selections are applied to the muon candidates. Muon candidates with an MDC track polar angle satisfying $0.82<|\!\cos\theta|<0.86$ are vetoed because of the gap of barrel and endcap of EMC. The track direction difference of the muon candidate in the MDC and MUC is used to constrain background. The polar angle difference between the track in the MUC and MDC is required to be less than $20^{\circ}$, and the azimuthal angle difference between the track in the MUC and MDC is required to be less than $30^{\circ}$. A detailed discussion can be found in Ref. [52].
@@
Since the final state electron and muon are back-to-back, the opening angles in the polar and azimuthal directions are required to be less than $1.4^{\circ}$ and $1.7^{\circ}$, respectively. The opening angle in the polar direction is defined as $|\Delta\theta|=|180^{\circ}-(\theta_{1}+\theta_{2})|$, where $\theta_{1}$ and $\theta_{2}$ are the corresponding polar angles of the electron and muon, respectively. The opening angle in the azimuthal direction is defined as $|\Delta\phi|=|180^{\circ}-|\phi_{1}-\phi_{2}||$, where $\phi_{1}$ and $\phi_{2}$ are the azimuthal angles of the electron and muon, respectively.
@@
To suppress background from the processes $e^{+}e^{-}\to\gamma e^{+}e^{-}$ and $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$, the number of detected photons is required to be zero The photon candidates used for this veto are reconstructed using showers in the EMC detector. The showers are required to have a deposited energy greater than $25~\rm MeV$ in the barrel region ($|\!\cos\theta|<0.80$) and $50~\rm MeV$ in the end-cap region ($0.86<|\!\cos\theta|<0.92$). The difference between the EMC time and the event start time for each photon must be within [0, 700] ns to suppress electronic noise and showers unrelated to the event. To exclude showers originating from the final state charged tracks, the opening angle between the position of each shower in the EMC and the closest extrapolated charged track must be greater than 10 degrees.
@@
The normalized momentum sum $|\sum_{i}\vec{p_{i}}|/\sqrt{s}$ and the normalized energy sum $\sum_{i}E_{i}/\sqrt{s}$ are used to determine the number of signal events in the full dataset, where $\vec{p}_{i}$ is the momentum and $E_{i}$ is the reconstructed energy of final state particle $i$, and $\sqrt{s}$ is the c.m. energy. The signal region is defined by $|\sum_{i}\vec{p_{i}}|/\sqrt{s}\leq 0.03$ and $0.97\leq\sum_{i}E_{i}/\sqrt{s}\leq 1.04$. The selection criteria discussed above are optimized by maximizing the figure-of-merit (FOM) suggested by Punzi [53], given by
@@
The detection efficiency for the signal process is estimated using the signal MC sample, which is determined to be $(24.18\pm 0.10)\%$. The distribution of the signal MC sample inside and outside the signal region are shown in Fig 3(a). After applying all the selections to the data from $2367.0\times 10^{6}$ $\psi(3686)$ events, eight $\psi(3686)\to e^{\pm}\mu^{\mp}$ candidates are observed in the signal region, as shown in Fig. 3(b).
4 Background study